# Blending 3D Geometry and Machine Learning for Multi-View Stereopsis

### v1

@@ 3.6 Cost Function
GC-MVSNet [17] 0.330 0.260 0.295 0.48 0.89 0.97
GC-MVSNet++ (ours) 0.319 0.246 0.2825 0.407 0.702 0.81
3.7 Depth Filtering
To mitigate the impact of inconsistent depth values, we apply the same GC module (refer to Alg. 1) to filter accurate depth maps, with a minor adjustment. Specifically, we substitute the estimated reference depth map ($D_{0}$) with its ground truth counterpart ($D_{0}^{gt}$) in Alg. 1. The hyperparameters are set to $D_{pixel}=2$, $D_{depth}=0.25$, and $M=8$, ensuring that only the most erroneous depth values are filtered out. Fig. 6 displays the original and filtered depth maps from various viewpoints of a sample scene in the DTU dataset. The left half of the figure illustrates viewpoints with the most filtered-out values, while the right half shows viewpoints with the fewest. Although we initially identified this depth map rendering issue in the DTU dataset, the approach is broadly applicable to other datasets employing similar methods for generating ground truth depth maps. For the BlendedMVS dataset, we use $D_{pixel}=0.5$, $D_{depth}=0.05$, and $M=10$ in the depth filtering process.
4 Experiments
4.1 Implementation Details
Following general practice [5], we first train and evaluate our model on DTU. Then, we finetune on BlendedMVS to evaluate on Tanks and Temples. For training on DTU, we set the number of input images to $N=5$ and image resolution as $512\times 640$. The depth hypotheses are sampled from $425mm$ to $935mm$ for coarse-to-fine regularization with the number of plane sweeping depth hypotheses for the three stages set to 48, 32, and 8. The corresponding depth interval ratio (DIR) is set as 2.0, 0.8, and 0.4. The model is trained with Adam [62] for 9 epochs with an initial learning rate ($LR_{DTU}$) of 0.001, which decays by a factor of 0.5 once after $8^{th}$ epoch. For the Geometric Consistency (GC) module, we use $M$=8 and set the stage-wise thresholds $D_{pixel}$ as 1, 0.5, 0.25 and ${D_{depth}}$ as 0.01, 0.005, 0.0025. We use $\alpha$=$\beta$=1 and $\gamma=2$ for all experiments. We train our model with a batch size of 2 on 8 NVIDIA RTX A6000 GPUs for about 9 hours.
@@ 4.1 Implementation Details
GC-MVSNet++ (ours) 66.28 82.72 71.05 58.18 65.40 65.63 64.95 61.47 60.87 42.14 27.74 47.84 40.24 52.32 37.48 47.23
Figure 7: Visual comparison of reconstructed point clouds of GC-MVSNet with CasMVSNet [10], R-MVSNet [8] and Ground truths. Our method obtains a more complete point cloud. See Appendix H in Supplemental Material for all point clouds.
@@ 4.2 Experimental Performance
Evaluation on DTU
On DTU, we generate depth maps with $N$=5 at an input resolution of $864\times 1152$. We slightly adjust the depth interval ratio (DIR) to $1.6,0.7,0.3$ to accommodate the resolution change (more on DIR in Appendix C in Supplemental Material) and use the Fusibile algorithm [1] for depth fusion. Table 1 shows quantitative evaluations, where accuracy is the mean absolute distance in $mm$ from the reconstructed point cloud to the ground truth point cloud, completeness measures the opposite (see Appendix E in Supplemental Material), and overall is the average of these metrics indicate the overall performance of the models. We find that GC-MVSNet++ achieves the best overall score as well as the best completeness score, when compared to nearly two dozen previous and state-of-the-art techniques. Snapshots of the DTU test set point clouds are shown in the Supplementary Materials. We find that our model generates denser and more complete point clouds.
Evaluation on BlendedMVS
Unlike DTU and Tanks and Temples, evaluation on Blended MVS is usually measured as the quality of depth maps, not the quality of point clouds. We set $N$=5, $M$=8, image resolution as $576\times 768$, number of depth planes $D$=128, and finetune for 12 epochs with one-tenth the learning rate we used for DTU ($\frac{1}{10}LR_{DTU}$). We follow [57] for the evaluation process.
Table 1 presents the results of our quantitative evaluation, using three metrics: Endpoint error (EPE) is the average $L_{1}$ distance between the estimated and the ground truth depth values, and $e_{1}$ and $e_{2}$ are the ratio of number of pixels with $L_{1}$ error larger than 1mm and 3mm, respectively. The significant improvement in depth map estimates corroborates that providing explicit geometric cues during training helps the model learn about multi-view geometric consistency while requiring many fewer training iterations. The addition of dense-CostRegNet and the use of the depth filtration module led to significant quantitative and qualitative improvements for GC-MVSNet++. See Appendix G of the Supplemental Material for point clouds.
@@
Evaluation on Tanks and Temples
We also test the performance of our model on an outdoor dataset with the Tanks and Temples benchmark. To adapt to this change, we first finetune our model on BlendedMVS and then evaluate on the intermediate and advanced test sets of Tanks and Temples. We use an image resolution of $576\times 768$, $N$=7, $M$=10, one-tenth the learning rate of DTU ($\frac{1}{10}LR_{DTU}$), and $D$=192 for finetuning. We finetune the model for 12 epochs. The camera parameters and neighboring view selection are used as in R-MVSNet [8] and follow evaluation steps described in MVSFormer [54]. Table 2 presents our quantitative comparison of different methods. GC-MVSNet++ achieves second best on the intermediate set and the best on the advanced set evaluation. Fig. 8 shows point clouds visualizing precision and recall comparisons for Train (intermediate set) and Temple (advanced set) with other MVS methods. These plots are downloaded from the Tanks and Temple benchmark leaderboard. See Appendix G in Supplemental Material for point clouds.
@@ 4.3 Ablation Study
Having demonstrated the efficacy of our approach relative to the state of the art, we now conduct ablation studies to evaluate the importance of the various components of our model.
Range of $\xi_{p}$
@@
$\xi_{p}$ is generated using the mask sum ($mask\_sum$ in Alg. 1), and is the sum of penalties accumulated across the $M$ source views during multi-view geometric consistency check. At this stage, its elements take a discrete value between 00 and $M$. Using mask sum as-is leads to a very high penalty per-pixel, and thus a very high loss value. Such a high loss value destabilizes the learning process. We control the magnitude of the penalty by controlling the range of the per-pixel penalty.
@@
We explore two different ranges, $[1,2]$ and $[1,3]$. To generate $\xi_{p}\in[1,2]$, we divide the mask sum by $M$ and then add $1$. To generate $\xi_{p}\in[1,3]$, we divide the mask by $\frac{M}{2}$ and then add $1$. Table 5 shows the impact of these two ranges for $M$=8. Since $\xi_{p}\in[1,2]$ produces the best results, we use it for all other experiments.
Hyperparameters of GC module
@@
The GC module has two types of hyperparameters, global and local. In this section, we investigate the effect of these hyperparameters on our results.
@@
The global hyperparameter $M$ is the number of source views across which the geometric consistency is checked, and is the same for all three stages (coarse, intermediate, and refinement stages). For training on DTU, we vary the value of $M$ while keeping $N$=5, i.e. while the MVS method uses only $4$ source views to estimate $D_{0}$, the GC module checks the geometrical consistency of $D_{0}$ across $M$ source views. It is important to note that the first $N$-1 out of $M$ source views are exactly the same used by GC-MVSNet++ to estimate $D_{0}$. We always keep $M\geq N-1$.
@@
Table 5 presents a quantitative comparison for different values of $M$ and the number of training iterations required for optimal performance of the model. At $M=N-1=4$, i.e. checking geometric consistency across the same number of source views as used by GC-MVSNet++ to estimate $D_{0}$, the model performance significantly improves with a sharp decrease in training iteration requirements, as compared to our baseline GC-MVS-base (Table 6). As we increase the value of $M$ from $4$ to $10$, the training iterations required by our model further decrease. We find that at $M=8$, which is twice the number of source views used by GC-MVSNet++, it achieves its best performance.
@@
The two local hyperparameters, $D_{pixel}$ and $D_{depth}$, are the stage-wise thresholds applied to generate $PDE_{mask}$ and $RDD_{mask}$ in Alg. 1. These values are set to smaller values in the later (finer) stages, providing a stricter penalty to geometrically inconsistent pixels at finer resolutions. Table 5 shows the overall performance of GC-MVSNet++ with a range of different $D_{pixel}$ and $D_{depth}$ thresholds. GC-MVSNet++ performance remains fairly consistent and achieves its best performance with $D_{pixel}=1,0.5,0.25$ and $D_{depth}=0.01,0.005,0.0025$. We use these threshold values for all datasets throughout the paper.
@@
Table 3: Impact of range of $\xi_{p}$ during training on DTU with $M$=8, $N$=5. Numbers are generated on DTU evaluation set.
@@
Table 5: Overall score on the evaluation set of DTU [18] for different values of $D_{depth}$ and $D_{pixel}$. $M$ is fixed at $8$. C, I, and R means Coarse, Intermediate, and Refine stages.
@@
Our Geometric Consistency module is generic and can be integrated into many different MVS pipelines [^1] We only argue about the GC module as a plug-in to improve other MVS methods, and do not include our proposed dense CostRegNet in any of these methods. Including it will change the original design of these methods.. To demonstrate this, we tested it with two very different MVS pipelines, TransMVSNet [5] and CasMVSNet [10]. CasMVSNet treats depth estimation as a regression problem, while TransMVSNet treats it as a classification problem and uses winner-take-all to estimate the final depth map. We purposefully choose different methods to show that the GC module can perform well for both types of formulation. We compare the architectures of GC-MVSNet++ with TransMVSNet and CasMVSNet in Sec. 5.
@@
Table 6 presents the results, showing the impact of adding the GC module, the deformable feature extraction network (shown as other modifications in the table), and depth filtration in the original method pipeline. To observe the absolute impact of adding these modifications, we do not modify the original pipelines in any other way. We do not include our proposed dense-CostRegNet in the original methods as it would enhance their learning capabilities. We focus on the extent of improvement in the original method with the integration of the GC module and depth filtration preprocessing step. We observe in the table that applying only the other modification leads to degradation in performance for both methods. It indicates that the other modification helps in stabilizing the training process and promoting reproducibility but has no significant impact on the performance of the model on its own. We also observe a sharp increase in model performance and a decrease in training iteration requirements after integrating our GC module into the original pipeline. With GC, training the CasMVSNet pipeline requires only $11$ epochs instead of $16$ epochs, while TransMVSNet (with GC module) requires only $8$ epochs instead of $16$ epochs. This corroborates our hypothesis that multi-view geometric consistency significantly reduces training computation because it accelerates the geometric cues learning. Applying the depth filtration preprocessing along with the GC module further improves the performance of both methods. Eliminating the erroneous ground truth pixels from the learning process through filtration provides more consistent geometric cues learning.
Stages of GC-MVSNet++
Table 6 also shows different stages of development of GC-MVSNet++. GC-MVS-base uses TransMVSNet pipeline with cross-entropy loss and performs much worse than original TransMVSNet [5] which uses focal loss. With only D-FEN modifications, it slightly improves the overall performance indicating better feature quality with deformable convolution. However, D-FEN does not reduce the training iteration requirements on its own. Only after applying the GC module, independently and with D-FEN, we see a significant reduction in training iteration requirements and a significant improvement in the overall accuracy metric. This clearly shows the significance of multi-view multi-scale geometric consistency checks in the GC-MVSNet++ pipeline. Including the DF preprocessing step further aids the GC module by eliminating erroneous ground truth values from the training loop and provides more consistent geometry cues across multiple source views. Finally, adding dense-CostRegNet with all other modifications improves the cost regularization process with its advanced architecture and dense identity connection. This leads to a significant jump in the overall performance of GC-MVSNet++.
@@
$L_{1}$checkmark $\times$ $\checkmark$ 0.330 11
@@
Table 7: Stages of performance improvement of GC-MVSNet++ with different modifications on DTU [18]. It uses Cross-entropy loss [13] for training. DF, D-FEN, GC, and DenseCostReg indicate depth filtering, Deformable-Feature Extraction Network, Geometric Consistency module, and Dense Cost Regularization network, respectively.
@@
GC-MVS-base $\times$ $\times$ $\times$ $\times$ 0.332
GC-MVS-base $\checkmark$ $\times$ $\times$ $\times$ 0.328
GC-MVS-base $\times$ $\checkmark$ $\times$ $\times$ 0.298
GC-MVSNet $\checkmark$ $\checkmark$ $\times$ $\times$ 0.295
GC-MVSNet $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 0.291
GC-MVSNet++ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.2825

### v2

@@ 3.7 Depth Filtering
MVSFormer++ [68] 0.309 0.252 0.2805
GoMVS [69] 0.347 0.227 0.287
GC-MVSNet [19] 0.330 0.260 0.295 0.48 7.48 2.78
GC-MVSNet++ (ours) 0.319 0.246 0.2825 0.407 5.79 2.41
3.8 GC-MVSNet++ Design Insights
The typical supervised MVS pipeline has three interconnected parts: feature extraction, cost volume regularization, and the supervision signal. The feature extraction networks generate initial features used to form the initial cost volume, which is then refined through cost volume regularization to estimate depth. The learning process is driven by the supervision signal provided by the loss function.
Our method begins by enhancing supervision with geometric consistency, explicitly penalizing depth estimates lacking consistency across multiple views. This geometry-aware supervision encourages improved feature extraction with better physical understanding of the 3D scenes and more accurate initial cost volume creation. However, the improved feature extraction and initial cost volume are limited by the rigidity of sampling locations in convolutional layers and the lack of global aggregation. To address this, we first add a global feature aggregation module for intra-image and inter-image feature fusion, and then adopt deformable convolution layers to allow flexible sampling that better accommodates geometric and view variations. (See Table 7 for ablation experiments.)
Inspired by the benefits of flexible feature sampling, we sought a parallel form of flexibility within the cost volume regularization network. Dense network connectivity, as exemplified by DenseNet [52], naturally supports flexible, multi-scale feature reuse with strong gradient flow for optimization [53, 52, 70], thereby enhancing learning representational power. Leveraging this insight, we adopted dense connectivity inspired modules for increased feature reuse, stronger gradient flow, and improved landscape for optimization [71, 70]. While the encoding module (simple dense block) closely follows a traditional DenseNet for its enhanced encoding capabilities [52], we designed a feature-dense decoding module specifically tailored for cost volume refinement. As illustrated in Fig. 4(f), it consists of multiple feature-dense block units, each containing a 3D convolution, group normalization, and ReLU activation. Outputs from these units are concatenated, progressively expanding the receptive field and facilitating accurate estimation across multiple depth planes. This compact yet effective design significantly improves cost volume regularization performance.
Together, these modules form a powerful pipeline comprising a flexible feature extraction network with global feature aggregation, and a robust cost regularization network featuring a densely connected encoder for rich representation learning and a feature-dense decoder with an expanding receptive field to effectively handle multiple depth planes estimates. Empirical validations underscore the efficacy of these components, collectively driving significant improvements in the MVS pipeline.
@@ 3.8 GC-MVSNet++ Design Insights
GC-MVSNet++ 66.28 82.72 71.05 58.18 65.40 65.63 64.95 61.47 60.87 42.14 27.74 47.84 40.24 52.32 37.48 47.23
4 Experiments
4.1 Implementation Details
Following general practice [5], we first train and evaluate our model on DTU. Then, we finetune on BlendedMVS to evaluate on Tanks and Temples. For training on DTU, we set the number of input images to $N=5$ and image resolution to $512\times 640$. The depth hypotheses are sampled from $425mm$ to $935mm$ for coarse-to-fine regularization with the number of plane sweeping depth hypotheses for the three stages set to 48, 32, and 8. The corresponding depth interval ratio (DIR) is set as 2.0, 0.8, and 0.4. The model is trained with Adam [72] for 9 epochs with an initial learning rate ($LR_{DTU}$) of 0.001, which decays by a factor of 0.5 once after $8^{th}$ epoch. For the GC-module, we use MM=8 and set the stage-wise thresholds $D_{pixel}$ as 1, 0.5, 0.25 and $D_{depth}$} as 0.01, 0.005, 0.0025. We use $\alpha$$\beta$=1 and $\gamma=2$ for all experiments. We train our model with a batch size of 2 on 8 NVIDIA RTX A6000 GPUs for about 9 hours.
@@ 4.2 Experimental Performance
Evaluation on DTU: We generate depth maps with NN=5 at an input resolution of $864\times 1152$. We slightly adjust the depth interval ratio (DIR) to $1.6,0.7,0.3$ to accommodate the resolution change (more on DIR in Appendix C) and use the Fusibile algorithm [1] for depth fusion. Table 1 shows quantitative evaluations, where accuracy is the mean absolute distance in $mm$ from the reconstructed point cloud to the ground truth point cloud, completeness measures the opposite (more details in Appendix E), and overall is the average of these metrics indicate the overall performance of the models. We find that GC-MVSNet++ achieves the second best overall and completeness scores when compared to previous state-of-the-art techniques. Snapshots of the DTU test set point clouds are shown in the Supplementary Materials. We find that our model generates denser and more complete point clouds.
Evaluation on BlendedMVS: Unlike DTU and Tanks and Temples, evaluation on Blended MVS is usually measured as the quality of depth maps, not the quality of point clouds. First, we finetune our model for 12 epochs with NN=5, MM=8, number of depth planes DD=128, at a resolution of $576\times 768$, with one-tenth the learning rate of DTU ($\frac{1}{10}LR_{DTU}$).
Following evalutation process of Darmon et al. [64], we generate Table 1 for a quantitative comparison with other methods using three metrics: Endpoint error (EPE) is the average $L_{1}$ distance between the estimated and the ground truth depth values, and $e_{1}$ and $e_{2}$ are the ratio of number of pixels with $L_{1}$ error larger than 1mm and 3mm, respectively. The significant jump in depth estimates corroborates our intuition that providing explicit geometric cues during training helps the model be mindful about preserving the geometric consistency of a view during inference. The addition of dense-CostRegNet and the use of the depth filtration module led to significant quantitative and qualitative improvements for GC-MVSNet++. error maps and point clouds are shown in Appendices H and I, respectively.).
@@
[…]
@@ 4.3 Ablation Study
We conduct ablation studies to evaluate the importance of the various components of GCMVSNet++. We provide detailed comparisons for inference time and memory requirements with MVSFormer++, TransMVSNet, CasMVSNet, and GC-MVSNet in Appendix F of the supplementary material.
@@
$\mathbf{\xi_{p}}$ Range: $\xi_{p}$ is generated using the mask sum (mask_summask\_sum in Alg. 1), and is the sum of penalties accumulated across the $M$ source views during multi-view geometric consistency check. At this stage, its elements take a discrete value between 0 and $M$. Using mask sum as-is leads to a very high penalty per-pixel, and thus a very high loss value, destabilizing the learning process. We control the magnitude of the penalty by controlling the range of the per-pixel penalty.
@@
We explore two different ranges, $[1,2]$ and $[1,3]$. To generate $\xi_{p}\in[1,2]$, we divide the mask sum by $M$ and then add $1$. To generate $\xi_{p}\in[1,3]$, we divide the mask by $\frac{M}{2}$ and then add $1$. Table 5 shows the impact of these two ranges for MM=8. Since $\xi_{p}\in[1,2]$ produces the best results, we use it for all other experiments.
@@
Hyperparameters of GC module The GC module has two types of hyperparameters, global and local. In this section, we investigate the effect of these hyperparameters on our results.
@@
The global hyperparameter $M$ is the number of source views across which the geometric consistency is checked, and is the same for all three stages (coarse, intermediate, and refinement stages). For training on DTU, we vary the value of $M$ while keeping NN=5, i.e. while the MVS method uses only $4$ source views to estimate $D_{0}$, the GC module checks the geometrical consistency of $D_{0}$ across $M$ source views. It is important to note that the first NN-1 out of $M$ source views are exactly the same used by GC-MVSNet++ to estimate $D_{0}$. We always keep $M\geq N-1$.
@@
Table 5 presents a quantitative comparison for different values of $M$ and the number of training iterations required for optimal performance of the model. At MM=N1N-1= 44, i.e. checking geometric consistency across the same number of source views as used by GC-MVSNet++ to estimate $D_{0}$, the model performance significantly improves with a sharp decrease in training iteration requirements, as compared to our baseline GC-MVS-base. As we increase the value of $M$ from $3$ to $10$, the training iterations required by our model further decrease. We find that at $M=8$, which is twice the number of source views used by GC-MVSNet++, it achieves its best performance.
@@
The two local hyperparameters, $D_{pixel}$ and $D_{depth}$, are the stage-wise thresholds applied to generate $PDE_{mask}$ and $RDD_{mask}$ in Alg.1. These are set to smaller values in the later (finer) stages, providing a stricter penalty to geometrically inconsistent pixels at finer resolutions. Table 5 shows the overall performance of GC-MVSNet++ with a range of different $D_{pixel}$ and $D_{depth}$ thresholds. GC-MVSNet++ performance remains fairly consistent and achieves its best performance with $D_{pixel}=1,0.5,0.25$ and $D_{depth}=0.01,0.005,0.0025$. We use these threshold values throughout the paper.
@@
Table 3: Impact of range of $\xi_{p}$ during training on DTU with MM=8, NN=5. Numbers are generated on DTU evaluation set.
@@
Table 5: Overall score on the evaluation set of DTU [21] for different values of $D_{depth}$ and $D_{pixel}$. $M$ is fixed at 88. C, I, and R means Coarse, Intermediate, and Refine stages.
@@
GC-module as a plug-in: Geometric Consistency module is generic and can be integrated into many different MVS pipelines To demonstrate this, we test it with two very different MVS pipelines, TransMVSNet and CasMVSNet. CasMVSNet treats depth estimation as a regression problem, while TransMVSNet treats it as a classification problem and uses winner-take-all to estimate the final depth map. We purposefully choose different methods to show that the GC module can perform well for both types of formulation. We compare the architectures of GC-MVSNet++ with TransMVSNet and CasMVSNet in discussion section. We also compare it with MVSFormer and MVSFormer++.
@@
Table 7 presents the results, showing the impact of adding the GC module, the deformable feature extraction network (shown as other modifications in the table), and depth filtration in the original method pipeline. To observe the absolute impact of adding these modifications, we do not modify the original pipelines in any other way. We do not include our proposed dense-CostRegNet or Deformable feature extraction network in the original methods as it would enhance their learning capabilities. We focus on the extent of improvement in the original method with the integration of the GC module and depth filtration preprocessing step. We observe in the table that applying only the other modification leads to degradation in performance for both methods. It indicates that the other modification helps in stabilizing the training process and promoting reproducibility but has no significant impact on the performance of the model on its own. We also observe a sharp increase in model performance and a decrease in training iteration requirements after integrating our GC module into the original pipeline. With GC, training the CasMVSNet pipeline requires only $11$ epochs instead of $16$ epochs, while TransMVSNet (with GC module) requires only $8$ epochs instead of $16$ epochs. This corroborates our hypothesis that multi-view geometric consistency significantly reduces training computation because it accelerates the geometric cues learning. Applying the depth filtration preprocessing along with the GC module further improves the performance of both methods. Eliminating the erroneous ground truth pixels from the learning process through filtration provides more consistent geometric cues learning.
@@
$L_{1}$ $\checkmark$ $\times$ $\checkmark$ 0.330 11
@@
Table 7: Stages of performance improvement of GC-MVSNet++ with different modifications on DTU [21]. It uses Cross-entropy loss [13] for training. DF, D-FEN, GC, GFA, and DenseCostReg indicate depth filtering, Deformable-Feature Extraction Network, Geometric Consistency module, Global Feature Aggregation and Dense Cost Regularization network, respectively.
@@
GC-MVS-Base $\times$ $\times$ $\times$ $\times$ $\times$ 0.332
GC-MVS-Base $\times$ $\checkmark$ $\times$ $\times$ $\times$ 0.334
GC-MVS-Base $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ 0.328
GC-MVS-Base $\times$ $\times$ $\checkmark$ $\times$ $\times$ 0.298
GC-MVS-Base $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 0.295
GC-MVS-Base $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 0.291
GC-MVSNet++ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.2825
Stages of GC-MVSNet++: Table 7 shows different stages of development of GC-MVSNet++. GC-MVS-base uses the TransMVSNet pipeline (without the feature matching transformer and adaptive receptive field modules) with cross-entropy loss and performs much worse than original TransMVSNet which uses focal loss. Including the global feature aggregation module with the basic feature extraction network performs roughly similar. But including D-FEN (Deformable-Feature Extraction Network) modifications slightly improves the overall performance indicating better feature quality with deformable convolution for global feature aggregation. However, D-FEN does not reduce the training iteration requirements on its own, see Table. 7. Only after applying the GC module, independently and with D-FEN, we see a significant reduction in training iteration requirements and a significant improvement in the overall accuracy metric. This clearly shows the significance of multi-view multi-scale geometric consistency checks in the GC-MVSNet++ pipeline. Including the DF preprocessing step further aids the GC module by eliminating erroneous ground truth values from the training loop and provides more robust geometry cues across multiple source views. Finally, adding dense-CostRegNet with all other modifications improves the cost regularization process with its advanced architecture and dense identity connection, leading to a significant improvement.